OpenStudy (blackbird02):

A fishing boat left the port and cruised for 32 miles at a heading of 24 degrees. It then cruised for 21 miles at a heading of 114 degrees. How far was the boat now located from the dock?

OpenStudy (nfcfox):

Here are some basic equations to solver this problem and here is a drawing of what I'm assuming is happening and hopefully you will agree. \[\cos(\theta)=\frac{ adjacent }{ hypotenuse }\] \[\sin(\theta)=\frac{ opposite }{ hypotenuse }\] 

OpenStudy (campbell_st):

well I think there are some problems with the diagram.. and bearing are measured from north in a clockwise direction  then you can use the law of cosines to solve th rpoblem